# Fashion Image-to-Image Translation for Complementary Item Retrieval

### v2

@@ 4.3. Generative Compatibility Model (GeCo)
The GeCo model can create meaningful queries by harnessing realistic templates generated by the CIGM model. In contrast to previous approaches, GeCo integrates the templates into the query (Liu et al., 2020a, b; Lin et al., 2019), thereby adhering to a CoIR task framework. As discussed before, this method assigns greater significance to the generated templates, rather than treating them as regularization signals. Consequently, we consider only the visual signal to evaluate the effectiveness of our methodology.
The objective of the learning procedure is to align the query and bottom embedding space better, thereby maximizing the compatibility between the fashion items. Specifically, GCGRM employs a pre-trained ResNet18 (He et al., 2016) backbone to project the triplet $(t_{i},\hat{b}_{i},b_{j})$ onto a feature space $F\subset\mathbb{R}^{512}$. Subsequently, the latent representations $\tilde{t_{i}},\tilde{b_{i}}$ are concatenated and projected via a linear layer onto a lower-dimensional space $W\subset\mathbb{R}^{128}$. Similarly, $\tilde{b_{j}}$ is projected onto another space $\tilde{W}\subset\mathbb{R}^{128}$, with the objective of aligning the spaces $W$ and $\tilde{W}$igure 6 depicts the overall structure of the GeCo model.
We argue that maximizing the compatibility is equivalent to maximize the similarity between vectors lying in the two spaces $W$ and $\tilde{W}$, measured via a simple dot product. To achieve this objective, we have formulated the loss function taking into account both BPR (Rendle et al., 2009) and Contrastive Learning objectives, namely the Information Noise-Contrastive Estimation (InfoNCE) (van den Oord et al., 2018) loss. While the BPR loss focuses on optimizing pairwise rankings by ensuring preferred items are ranked higher than non-preferred items, the InfoNCE loss operates in a self-supervised manner, aiming to bring positive sample pairs closer together in the latent space while pushing negative sample pairs farther apart. Thus, optimizing these two loss functions leads to maximizing the similarity among compatible pairs while considering all the other bottoms as negatives.
Formally, given a set $\mathcal{Q}\coloneqq\{(i,j,k)\mid(t_{i},b_{j})\in\mathcal{P},\quad b_{k}\in\mathcal{P}\setminus b_{j}\}$, we define the compatibility score between the top-positive pair and the top-negative ones in the Equation 9, where $m_{ij}$ represents the positive score, $m_{ik}$ represents the negative scores, and $\text{sim}(\cdot,\cdot)$ denotes the dot product.
@@
We define the BPR loss in Equation 11 and the InfoNCE loss in Equation 12, where $\tau$ is the temperature parameter, and the regularization component in Equation 13 with $\theta_{c}$ representing the parameters of the GCGRM model.
@@ 5.1. Datasets
columneven = c, column3 = c, hline1-2,5 = -, Dataset & Tops Bottoms Positives
FashionVC (Song et al., 2017) 12,309 11,281 17,140
ExpReduced (Lin et al., 2020) 11,291 9,575 17,140
FashionTaobaoTB (ours) 32,394 25,687 76,426
@@ 5.2. Evaluation metrics
We compute the AUC by evaluating each top $t_{i}$ with its corresponding positive bottom $b_{j}$ and negative bottom $b_{k}$. It is noteworthy that in a prior work on compatibility modeling, Song et al. (2017) utilized 3 random negatives for AUC and 9 random negatives for MRR. Unfortunately, there is no clear reference for the methodology by which these negatives were sampled or justifying this choice. In contrast, we compute MRR by considering the rank position of the positive bottom $b_{j}$ within the ranked list, where all other bottoms in the test set are treated as negatives.
@@ 5.4. Baseline Models
  • BPR-DAE (Song et al., 2017): BPR-DAE is a content-based clothing matching scheme that jointly models the compatible preferences of fashion items with multi-modalities and the coherent relation among different modalities of the same item. We adapt this method to encode only the visual features using the original AlexNet backbone, as discussed and evaluated in (Liu et al., 2020a).
  • MGCM (Liu et al., 2020a): MGCM is a compatibility schema based on a GAN model that computes fashion item compatibility by considering both item-item and item-template compatibility scores. The item-item score represents the actual compatibility between the top and the candidate bottom embeddings, while the item-template score is the $L_{1}$ norm between the candidate bottom and the generated embeddings.
@@
row1 = c, column3 = c, column5 = c, column7 = c, cell12 = c=2, cell14 = c=2, cell16 = c=2, cell22 = c, cell24 = c, cell26 = c, cell32 = c, cell34 = c, cell36 = c, cell42 = c, cell44 = c, cell46 = c, cell52 = c, cell54 = c, cell56 = c, cell62 = c, cell64 = c, cell66 = c, cell72 = c, cell74 = c, cell76 = c, hline1,3,8 = -, hline2 = 2-7, Model & FashionVC ExpReduced FashionTaobaoTB
@@
Table 2 presents the results obtained using the proposed evaluation protocol. It is evident that GeCo outperforms all other models across all datasets. Notably, BPR-DAE emerges as the second-best performer, exhibiting a significant margin over other models on the FashionVC and ExpReduced datasets. Conversely, MGCM achieves AUC values comparable to BPR-DAE only on the FashionTaobaoTB dataset, albeit still lower than GeCo, while exhibiting the second-highest MRR. This observation suggests that training an end-to-end generative compatibility model like MGCM yields high performance primarily when more information is available in the training data. Conversely, on smaller datasets such as FashionVC and ExpReduced, only simpler strategies like BPR-DAE demonstrate competitive performance. Other generative models struggle to achieve satisfactory results, thereby highlighting the efficacy of our model on these datasets.
We note that the baseline results reported in Table 2 differ from the results in (Liu et al., 2020a) due to differences in how the metrics are computed. As mentioned above, Liu et al. (2020a) calculated the AUC and MRR metrics in an uncommon approach by pairing each positive item with three and nine negative items, respectively. To understand how our proposed model would fare in such a setup, we repeated the measurement using the specific MRR and AUC calculation from (Liu et al., 2020a), leading to the results shown in Table 3. We find that also under this uncommon configuration, GeCo outperforms all considered baselines. BPR-DAE emerges as the second-best model in most cases, with the exception of the FashionTaobaoTB dataset where MGCM benefits from a larger volume of data, enabling it to learn more meaningful relations. Once more, our model demonstrates its adaptability to varying sample sizes in the training set.
@@
row1 = c, column3 = c, column5 = c, column7 = c, cell12 = c=2, cell14 = c=2, cell16 = c=2, cell22 = c, cell24 = c, cell26 = c, cell32 = c, cell34 = c, cell36 = c, cell42 = c, cell44 = c, cell46 = c, cell52 = c, cell54 = c, cell56 = c, cell62 = c, cell64 = c, cell66 = c, cell72 = c, cell74 = c, cell76 = c, hline1,3,8 = -, hline2 = 2-7, Model & FashionVC ExpReduced FashionTaobaoTB
@@ 5.5.1. Weight Variation Analysis
Overall, this analysis underscores the necessity of finely tuned hyperparameters, particularly the interplay between the BPR and InfoNCE components, which significantly affect both tasks. Additionally, the $\tau$ parameter is pivotal in achieving optimal performance. The trends shown in Figure 7 were consistently observed across other datasets, supporting our conclusions. In particular, for the FashionVC dataset, the best results are achieved with $\alpha=0.25$, $\beta=0.25$, and $\tau=0.5$, while for the ExpReduced dataset, the optimal combination is $\alpha=0.25$, $\beta=0.25$, and $\tau=0.5$. This analysis primarily focuses on the combination of loss parameters that optimize compatibility modeling, as measured by the AUC. Despite this, it is certainly possible to identify other combinations of these parameters that can push the ranking performance of our model at the expense of accuracy performance (e.g., as is the case with $\tau$), but this exploration and optimization is outside the scope of our work.
@@ 5.6. Comparison of Templates
Furthermore, we assess the impact of the randomly selected templates in the complementary item retrieval task, with results shown in Figure 9. The CIGM model generates templates of significantly higher quality compared to baseline methods, achieving a four-fold improvement in resolution. As highlighted by the red boxes in the figure, our model consistently outperforms the baselines by correctly identifying the positive bottom in the ranked list across all evaluated cases. It is noteworthy that integrating these templates into the retrieval pipeline sometimes leads to generated details that differ from those of the positive items while still preserving compatibility information. For instance, despite the visual differences in some generated template, the retrieved items remain visually compatible with the input top. This demonstrates the capability of our approach to enhance diversity in complementary retrieval results.
@@ 6. Conclusions
In this work, we aim to advance the field of fashion compatibility modeling and fashion image retrieval by proposing a novel compatibility schema, named GeCo, that integrates realistic generations, thereby enhancing the retrieval process. GeCo represents the first instance of a composed retrieval pipeline that utilizes generated items as conditioning signals in a two-stage approach, demonstrating increased reliability and stability, even with limited training data.
We conducted an extensive benchmarking activity with three well-known baselines for this task (i.e., BPR-DAE, MGCM, and Pix2PixCM), adopting two historical datasets for this domain and a new one proposed in our work, which is based on the data previously collected from the Taobao platform. We make this new dataset, named FashionTaobaoTB, publicly accessible to provide a new resource for this domain and to ensure better reproducibility of the work.
Our experiments demonstrate the effectiveness of our two-stage approach across all tested datasets for both compatibility modeling and complementary item retrieval tasks. Additionally, these experiments validate the proposed loss function, with an extensive discussion on the impact of its components provided in the corresponding section. Furthermore, unlike existing baselines, our approach generates realistic templates while effectively addressing the aforementioned tasks.
Future developments may focus on introducing textual signals to enhance retrieval performance. Investigations are required to determine the optimal integration point for these signals, whether at the GAN level or during retrieval. Additionally, incorporating personalization features and identifying the most effective stage for their integration into the retrieval pipeline represent promising avenues for future research.

### v3

@@ 4.3. Generative Compatibility Model (GeCo)
The GeCo model can create meaningful queries by harnessing realistic templates generated by the CIGM model. In contrast to previous approaches, GeCo integrates the templates into the query (Liu et al., 2020a, b; Lin et al., 2019), thereby adhering to a CoIR task framework. As discussed before, this method puts more emphasis to the generated templates, rather than treating them merely as regularization signals. Consequently, we consider only the visual signal to evaluate the effectiveness of our methodology.
The objective of the learning procedure is to better align the query and bottom embedding space better, thereby maximizing the compatibility between the fashion items. Specifically, GeCo employs a pre-trained ResNet18 (He et al., 2016) backbone to project the triplet $(t_{i},\hat{b}_{i},b_{j})$ onto a feature space $F\subset\mathbb{R}^{512}$. Subsequently, the latent representations $\tilde{t_{i}},\tilde{b_{i}}$ are concatenated and projected via a linear layer onto a lower-dimensional space $W\subset\mathbb{R}^{128}$. Similarly, $\tilde{b_{j}}$ is projected onto another space $\tilde{W}\subset\mathbb{R}^{128}$, with the objective of aligning the spaces $W$ and $\tilde{W}$igure 6 shows the overall structure of the GeCo model.
We argue that maximizing the compatibility is equivalent to maximize the similarity between vectors lying in the two spaces $W$ and $\tilde{W}$, measured via a simple dot product. To achieve this objective, we have formulated the loss function taking into account both BPR (Rendle et al., 2009) and Contrastive Learning objectives, namely the Information Noise-Contrastive Estimation (InfoNCE) (van den Oord et al., 2018) loss. While the BPR loss focuses on optimizing pairwise rankings by ensuring preferred items are ranked higher than non-preferred items, the InfoNCE loss operates in a self-supervised manner, aiming to bring positive sample pairs closer together in the latent space while pushing negative sample pairs farther apart. Thus, optimizing these two loss functions maximizes the similarity among compatible pairs while considering all the other bottoms as negatives.
Formally, given a set $\mathcal{Q}\coloneqq\{(i,j,k)\mid(t_{i},b_{j})\in\mathcal{P},\quad b_{k}\in\mathcal{P}\setminus b_{j}\}$, we define the compatibility score between the positive top-bottom pairs and the negative ones in the Equation 9, where $m_{ij}$ represents the positive score, $m_{ik}$ represents the negative scores, and $\text{sim}(\cdot,\cdot)$ denotes the dot product.
@@
We define the BPR loss in Equation 11 and the InfoNCE loss in Equation 12, where $\tau$ is the temperature parameter, and the regularization component in Equation 13 with $\theta_{c}$ representing the parameters of the GeCo model.
@@ 5.1. Datasets
cell22 = c, cell23 = c, cell24 = c, cell32 = c, cell33 = c, cell34 = c, cell42 = c, cell43 = c, cell44 = c, hline1-2,5 = -, Dataset & Tops Bottoms Positives
FashionVC  13,809 12,756 18,640
ExpReduced  12,791 11,026 18,640
FashionTaobaoTB (ours) 43,096 35,414 89,951
@@ 5.2. Evaluation metrics
We compute the AUC by evaluating each top $t_{i}$ with its corresponding positive bottom $b_{j}$ and negative bottom $b_{k}$. It is noteworthy that in a prior work on compatibility modeling, Song et al. (2017) utilized three random negatives for AUC and nine random negatives for MRR. Unfortunately, there is no clear reference for the methodology by which these negatives were sampled or justifying this choice, which is why we compute MRR by considering the rank position of the positive bottom $b_{j}$ within the ranked list, where all other bottoms in the test set are treated as negatives.
@@ 5.4. Baseline Models
  • Popularity: the compatibility score between top and bottom is based on “popularity”, quantified as the number of tops associated with the bottom $b_{j}$, following (Song et al., 2017).
  • BPR-DAE (Song et al., 2017): a content-based clothing matching scheme that jointly models the compatible preferences of fashion items with multi-modalities and the coherent relation among different modalities of the same item. We adapt this method to encode only the visual features using the original AlexNet backbone, as discussed and evaluated in (Liu et al., 2020a).
  • MGCM (Liu et al., 2020a): a compatibility schema based on a GAN model that computes fashion item compatibility by considering both item-item and item-template compatibility scores. The item-item score represents the actual compatibility between the top and the candidate bottom embeddings, while the item-template score is the $L_{1}$ norm between the candidate bottom and the generated embeddings.
@@
row1 = c, column3 = c, column5 = c, column7 = c, cell12 = c=2, cell14 = c=2, cell16 = c=2, cell22 = c, cell24 = c, cell26 = c, cell32 = c, cell34 = c, cell36 = c, cell42 = c, cell44 = c, cell46 = c, cell52 = c, cell54 = c, cell56 = c, cell62 = c, cell64 = c, cell66 = c, cell72 = c, cell74 = c, cell76 = c, cell82 = c, cell84 = c, cell86 = c, hline1,3,9 = -, hline2 = 2-7, Model & FashionVC ExpReduced FashionTaobaoTB
@@
Table 2 presents the results obtained using the proposed evaluation protocol. It is evident that GeCo outperforms all other models across all datasets. Notably, BPR-DAE emerges as the second-best performer, exhibiting a significant margin over other models on the FashionVC and ExpReduced datasets. Conversely, MGCM achieves AUC values comparable to BPR-DAE only on the FashionTaobaoTB dataset, albeit still lower than GeCo, while exhibiting the second-highest MRR. This observation suggests that training an end-to-end generative compatibility model, such as MGCM, delivers high performance mainly when the training data contains more information. The same holds for the other generative baseline used, namely Pix2PixCM. Conversely, on smaller datasets such as FashionVC and ExpReduced, only simpler strategies like BPR-DAE demonstrate competitive performance. Other generative models struggle to achieve satisfactory results, thereby highlighting the efficacy of our model on these datasets. Considering Popularity and Random baselines, we observe a not common behavior. Indeed, these results reveal that Random performs superior to Popularity: this is related to the distributions of pairs in the datasets analyzed. As shown in Figure 7, most bottoms are paired with only a limited number of tops, indicating that these bottoms are not really popular, as they tend to be recommended without a clear underlying rationale. Consequently, the Popularity baseline assigns relatively uniform compatibility scores across numerous items, which reduces its effectiveness and affects its performance in comparison to the Random baseline.
We note that the baseline results reported in Table 2 differ from the results in (Liu et al., 2020a) due to differences in how the metrics are computed. As mentioned above, Liu et al. (2020a) calculated the AUC and MRR metrics in an uncommon approach by pairing each positive item with three and nine negative items, respectively. To understand how our proposed model would fare in such a setup, we repeated the measurement using the specific MRR and AUC calculation from (Liu et al., 2020a), leading to the results shown in Table 3. We also find that  GeCo outperforms all considered baselines under this uncommon configuration. BPR-DAE emerges as the second-best model in most cases, except for the FashionTaobaoTB dataset, where MGCM benefits from a larger volume of data, enabling it to learn more meaningful relations. Once more, our model demonstrates its adaptability to varying sample sizes in the training set. Furthermore, Popularity and Random retain the same behavior as the previous evaluation setting.
@@
row1 = c, column3 = c, column5 = c, column7 = c, cell12 = c=2, cell14 = c=2, cell16 = c=2, cell22 = c, cell24 = c, cell26 = c, cell32 = c, cell34 = c, cell36 = c, cell42 = c, cell44 = c, cell46 = c, cell52 = c, cell54 = c, cell56 = c, cell62 = c, cell64 = c, cell66 = c, cell72 = c, cell74 = c, cell76 = c, cell82 = c, cell84 = c, cell86 = c, hline1,3,9 = -, hline2 = 2-7, Model & FashionVC ExpReduced FashionTaobaoTB
@@
22224444666688881010101012121212141414141616161618181818202020202222222224242424002,00020002{,}0002 , $0004{,}$ $0006{,}$ $0008{,}$  [^0] Number of paired topsFrequency
(a) FashionVC
22224444666688881010101012121212141414141616161618181818202020202222222224242424001,00010001{,}0001 , $0002{,}$ $0003{,}$ $0004{,}$ $0005{,}$  [^0] Number of paired topsFrequency
(b) ExpReduced
222244446666888810101010121212121414141416161616181818182020202022222222242424242,00020002{,}0002 , $0006{,}$ $00010{,}$ $00014{,}$  [^0] Number of paired topsFrequency
(c) FashionTaobaoTB
Figure 7. Distribution of pairs across all datasets.
@@ 5.5.1. Weight Variation Analysis
Overall, this analysis underscores the necessity of finely tuned hyperparameters, particularly regarding the interplay between the BPR and InfoNCE components, which significantly affect both tasks. Additionally, the $\tau$ parameter is pivotal in achieving optimal performance. The trends shown in Figure 8 were consistently observed across other datasets, supporting our conclusions. In particular, for the FashionVC dataset, the best results are achieved with $\alpha=0.25$, $\beta=0.25$, and $\tau=0.5$, while for the ExpReduced dataset, the optimal combination is $\alpha=0.25$, $\beta=0.25$, and $\tau=0.5$. This analysis primarily focuses on the combination of loss parameters that optimize compatibility modeling, as measured by the AUC. Despite this, it is certainly possible to identify other combinations of these parameters that can push the ranking performance of our model at the expense of accuracy performance (e.g., as is the case with $\tau$), but this exploration and optimization is outside the scope of our work.
@@ 5.6. Comparison of Templates
Furthermore, we assess the impact of the randomly selected templates in the complementary item retrieval task, with results shown in Figure 10. The CIGM model generates templates of markedly higher quality compared to baseline methods, ensuring a four-fold resolution improvement. As highlighted by the red boxes in the figure, our model consistently outperforms the baselines by correctly identifying the positive bottom in the ranked list across all evaluated cases. Notably, integrating these templates into the retrieval pipeline sometimes leads to generated details that differ from those of the positive items while preserving compatibility information. For instance, despite the visual differences in some generated templates, the retrieved items remain visually compatible with the input top. This demonstrates the capability of our approach to enhance diversity in complementary retrieval results.
@@ 6. Conclusions
In this study, we have advanced the field of fashion compatibility modeling and fashion image retrieval by introducing a novel compatibility schema, GeCo. This schema integrates realistic generations into both the compatibility modeling and retrieval processes, thereby enhancing the overall effectiveness of these tasks. To the best of our knowledge, GeCo is the first to introduce a composed retrieval pipeline that leverages generated items as conditioning signals in a two-stage approach. This innovative method has demonstrated marked improvements in reliability and stability, particularly in scenarios characterized by limited training data.
Notably, GeCo not only pioneers a two-stage approach but also introduces a new framework that redefines the compatibility modeling and top-bottom retrieval tasks as a composed image retrieval problem. Unlike traditional methods that rely on an image-text query pair, our model is the first to utilize a composed query comprising two images (e.g., a top image and a template image). Furthermore, GeCo extends the concept of composed image retrieval across entirely different domains (e.g., top and bottom distributions), as opposed to previous works that focus on retrieving items with minor modifications based on a textual query. A remarkable outcome of our study is the ability of GeCo to generate realistic templates that effectively address key challenges in these tasks, a capability not observed in existing baselines. The experimental results also reveal a critical gap in the literature concerning applying generative models in contexts with limited data.
Looking forward, several promising directions for future research emerge. One potential approach is the integration of textual signals within the developed framework to enhance retrieval performance further. Careful investigation will be needed to select the optimal point for incorporating these signals — whether at the GAN level or during the second stage of the GeCo. Additionally, exploring the integration of personalization features into the retrieval pipeline presents another promising area of research, to identify the most effective phase for their inclusion. These enhancements have the potential to further elevate the performance of fashion compatibility modeling and retrieval systems, providing users with more tailored and contextually relevant results.